OpenStudy (anonymous):

Find the value of the annuity a1=1000,I=.08,n= 10

OpenStudy (wolf1728):

I'm assuming it would be calculated as Total = Principal * (1 + rate) ^ years Total = 1,000 * (1 + .08) ^ 10 Total = 1,000 * 2.1589249973 Total = 2,158.92

OpenStudy (anonymous):

Just entered, it was wrong, correct answer was $14.486.56. I was stuck on that problem for more than an hour.

OpenStudy (wolf1728):

For 1,000 to mature into $14.486.56 would take a MUCH higher interest rate or a MUCH longer time.